OpenStudy (anonymous):

Simplify: i11√-121 Question options: 11 -11 11i -11i

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@dtan5457 the answer 11 was wrong sir :(

OpenStudy (anonymous):

and some others was wrong :(

OpenStudy (danjs):

\[\sqrt{-112} = \sqrt{-1 * 121} = \sqrt{-1}*\sqrt{121}\]

OpenStudy (danjs):

can you retype the original expression please

OpenStudy (anonymous):

i^11 sqaure root -121

OpenStudy (danjs):

oh i^11 power ok

OpenStudy (danjs):

\[i^11 = i^1*i^2*i^2*i^2*i^2*i^2\]

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (danjs):

i to the 11th power is (-1)(-1)(-1)(-1)(-1)*i

OpenStudy (dtan5457):

@EuniceOB97 sorry, but i dont see what i did wrong.

OpenStudy (danjs):

\[i ^{11} = -i\]

OpenStudy (anonymous):

-11i ?

OpenStudy (danjs):

\[\sqrt{-121} = \sqrt{-1}\sqrt{121} = 11i\]

OpenStudy (danjs):

11i * -i = -11i^2 = -11*(-1) = 11

OpenStudy (anonymous):

i=sqr(-1) i^2=-1 i^4=1

OpenStudy (anonymous):

they say 11 is wrong @DanJS

OpenStudy (dtan5457):

i dont see why eunice.

OpenStudy (anonymous):

i^11+i^1 =i^12=1

OpenStudy (danjs):

It is 11, i just checked on the calculator.

OpenStudy (anonymous):

121=11 so the answer =11

OpenStudy (danjs):

\[i ^{11}\sqrt{-121} = 11\]

OpenStudy (dtan5457):

@EuniceOB97 what others did i do wrong?

OpenStudy (dtan5457):

please post here so i can find mistakes

OpenStudy (dtan5457):

for this one i still got 11 so..

OpenStudy (anonymous):

Simplify: i10√-36 Question options: 6 -6 6i -6i

OpenStudy (dtan5457):

i^10 everybody

OpenStudy (anonymous):

Simplify: -i8√-400 Question options: 20 -20 20i -20i

OpenStudy (anonymous):

Simplify: | −11 + i | Question options: √122 12 11 + i 11 - i

OpenStudy (danjs):

i to the tenth times square root of -36 is -6i

OpenStudy (dtan5457):

i got -6i too..

OpenStudy (anonymous):

@dtan5457 do you see it ?

OpenStudy (anonymous):

the one with the absolute value i do not understand how it is wrong

OpenStudy (dtan5457):

im still sticking with -20i

OpenStudy (dtan5457):

for the 2nd q

OpenStudy (dtan5457):

honestly i dont see a single mistake in my work. maybe your answer key is messed up.

OpenStudy (anonymous):

do you guys know what the absolute value one is ?

OpenStudy (dtan5457):

got nothing but my old explanation^

OpenStudy (anonymous):

i^10*sqrt(-36)=i^10*i^1*sqrt(36)=i^(10+1)*6=I^11 6=-i 6

OpenStudy (anonymous):

-i^8*sqrt(-400)=-(i^(8+1))*20=-i^9 20=-(i) 20 = -i 20 =-20 i